What are semicarbaozones and semicarbazides?
What is the product when acetaldehyde reacts with Semicarbazides?

Grade:12th pass

1 Answers

Jayant Kishore
8 years ago
Dear Yogiraj,
H2N-NH-(C=O)-NH2 is semicarbazide.
H2*N-NH-(C=O)-NH2 + HCHO → H2C=*N-NH-(C=O)-NH
The Product formed is semicarbazone. This reaction is similar to the addition of ammonia derivatives to aldehydes and ketones . MOST IMPORTANT THING IS THAT ADDITION TAKES PLACE FROM NITROGEN ATOM (marked * here) WHOSE LONE PAIR ARE NOT IN RESONANCE WITH C=O IN SEMICARBAZIDE.